HIGH RESOLUTION SPECTROSCOPY OF NO IN HE-DROPLETS

A. METZELTHIN, ST. RUDOLPH,, K. VON HAEFTEN, M. HAVENITH, Physical Chemistry II, Ruhr-University Bochum, 44780 Bochum, Germany.

We have investigated the coupling of collective excitations in helium droplets to the rotation of light rotors. If we use NO as a molecular dopant it is possible to suppress this coupling effectively thereby obtaing high resolution spectra of NO in helium droplets. We have been able to resolve the Λ- doubling and the hyperfine structure of the Q(0.5) Π_{\frac{1}{2}} transition. Whereas the hyperfine structure remains unchanged compared to the gas phase, the Λ doubling is found to be considerably increased.